# Evidential Trust-Aware Model Personalization in Decentralized Federated Learning for Wearable IoT

### v1

@@ I Introduction
Wearable Internet of Things (IoT) devices such as smartwatches, fitness trackers, and medical monitors continuously collect sensor data for applications ranging from human activity recognition to early disease detection [chen2020fedhealth]. Training machine learning models on such data traditionally requires centralized collection, raising privacy concerns and imposing substantial communication and compute overheads on bandwidth-constrained edge networks.
Federated Learning (FL) addresses these concerns by enabling collaborative model training while keeping data on user devices [mcmahan2017communication]. However, standard centralized FL approaches create a single point of failure and scalability bottleneck problematic for IoT deployments spanning thousands of heterogeneous devices. Decentralized Federated Learning (DFL) eliminates this bottleneck through peer-to-peer architectures where devices exchange and aggregate model updates directly [lalitha2018fully], improving fault tolerance and reducing communication costs.
@@ I-B The Peer Compatibility Problem
Existing personalized FL methods typically assume centralized coordination where a server decides how to cluster or weight clients [tan2022towards]. In fully decentralized settings, each node must make local decisions about peer compatibility without central oversight.
@@ I-C Our Approach: From Uncertainty to Trust
We address this by leveraging uncertainty quantification. When a model evaluates data, we need to separate two fundamentally different types of uncertainty: uncertainty due to lack of knowledge about the data distribution (which signals incompatibility), and uncertainty due to inherent ambiguity in the data itself (which does not). Recent advances in evidential deep learning (EDL) [sensoy2018evidential] enable exactly this decomposition by having neural networks output not just predictions, but also quantified measures of these two uncertainty types.
@@ I-D Contributions
  • We introduce the first framework combining evidential uncertainty quantification with trust-aware personalization in DFL, enabling nodes to maintain personalized models through selective peer collaboration in non-IID environments.
@@ II-A Decentralized Federated Learning
FL was introduced to enable collaborative model training while keeping data on local devices [mcmahan2017communication]. The original approach uses a central server to coordinate training rounds: clients train locally, send updates to the server, which aggregates them and broadcasts the result. While this protects data privacy, it creates centralization challenges.
Decentralized FL removes the need for a central server entirely. Instead, nodes form a network and exchange model updates directly with their neighbors. This peer-to-peer approach improves fault tolerance and can reduce communication costs by leveraging local network structure. Researchers have established both theoretical convergence guarantees [lalitha2018fully, koloskova2020unified] and practical benefits including faster training on bandwidth-limited networks [lian2017can]. For IoT deployments where devices may join and leave dynamically, gossip-based communication protocols provide particularly natural patterns for information exchange [hegedHus2019gossip].
II-B Personalized Federated Learning
When devices have heterogeneous data distributions, forcing everyone to share a single global model often produces mediocre results. Personalized FL addresses this by letting each device maintain a model adapted to its own data while still benefiting from collaboration.
Early approaches include simple strategies like fine-tuning a global model on local data after training [wang2019federated], or treating each client as a related but distinct task in a multi-task learning framework [smith2017federated]. More sophisticated methods have emerged with theoretical guarantees. For example, some approaches maintain both a global model (learned collaboratively) and a personalized model (adapted locally), balancing the two with careful regularization [t2020personalized, li2021ditto]. Others apply meta-learning principles to learn models that can quickly adapt to new data distributions [fallah2020personalized].
Most personalized FL methods assume centralized coordination where a server decides how to group similar clients or weight their contributions. In decentralized settings without a server, personalization becomes harder. Recent work has explored peer selection mechanisms where nodes choose which neighbors to learn from based on task similarity [dai2022dispfl, fan2025pfeddst]. However, these approaches rely on heuristics such as cosine similarity of gradients or accuracy on shared test samples to assess peer compatibility rather than principled uncertainty quantification. These heuristics cannot distinguish between a peer with incompatible data and one with compatible but challenging data, limiting their effectiveness in highly heterogeneous environments.
@@ II-C Evidential Deep Learning
Standard neural network classifiers output probabilities for each class but cannot express confidence in these predictions. EDL [sensoy2018evidential] addresses this by having networks output evidence values that parameterize a Dirichlet distribution over probability vectors. This enables explicit quantification of epistemic uncertainty (lack of knowledge due to insufficient training or distribution shift) and aleatoric uncertainty (inherent data ambiguity). Epistemic uncertainty directly indicates whether a model has learned relevant patterns for given data, making it particularly useful for identifying compatible peers in heterogeneous FL settings. While EDL has been used in various FL contexts such as selecting reliable models [zhou2023trustworthy] or weighting aggregation in centralized settings [wang2023federated], no prior work has leveraged it for compatibility-based peer selection in fully decentralized personalized FL under non-IID data conditions.
@@ II-D Positioning and Research Gap
Table I compares Murmura with related approaches across key dimensions. As the table shows, no existing work combines all four capabilities needed for our setting: decentralized operation (no server), personalization (adapted models per node), principled uncertainty quantification for compatibility assessment, and trust-aware peer selection based on distributional compatibility.
@@
pFedMe [t2020personalized]
Ditto [li2021ditto]
DisPFL [dai2022dispfl]
PFedDST [fan2025pfeddst] Heuristic
D-PSGD [lian2017can]
FedUAA [wang2023federated] Central
TPFL [chen2024tpfl] Prediction-only
Murmura
@@ III-D Design Objectives
We design Murmura to simultaneously satisfy four objectives:
@@ IV Murmura Framework
This section presents the Murmura framework design, emphasizing the architectural decisions that enable systematic evaluation of trust-aware personalization under controlled heterogeneity.
@@ IV-A Design Principles and Architecture
Murmura achieves these goals through modular separation of concerns, as illustrated in Figure 1. The architecture comprises three layers. The Core layer separates network orchestration (managing training rounds and communication) from node-level decision making (local training and autonomous aggregation). This separation is essential: it allows individual nodes to implement different aggregation strategies while maintaining identical communication patterns, isolating the impact of personalization mechanisms from topology effects.
Figure 1: Murmura framework architecture showing the three-layer design. The Configuration Layer provides YAML/JSON-based experiment specification. The Core Components layer contains the Network Orchestrator and Node (Local Learner), along with Topology and Model definitions. The Pluggable Modules layer enables flexible integration of aggregation strategies and data partitioning schemes.
The Pluggable Modules layer decouples topology generation from aggregation logic. This enables systematic study of how personalization strategies perform under different network structures (ring, fully-connected, Erdős-Rényi [erdos1960evolution], k-regular) without conflating topology and algorithm effects—a common limitation in prior work where algorithms are evaluated on fixed topologies.
@@ V Evidential Trust-Aware Personalization
This section explains how Murmura uses evidential deep learning to assess peer compatibility and enable personalized model aggregation in heterogeneous data environments.
@@ V-B Training Evidential Models
Evidential models require a specialized loss function that encourages accurate predictions while preventing the model from simply claiming high confidence everywhere. Following Sensoy et al. [sensoy2018evidential], we use:
@@
where $\hat{p}_{k}=\alpha_{k}/S$ is the expected probability for class $k$, and the KL term prevents accumulating evidence for incorrect classes. The coefficient $\lambda_{t}$ gradually increases during training to allow initial learning before regularizing uncertainty.
@@ V-C Computing Compatibility from Uncertainty
2: Compute mean epistemic uncertainty $\bar{u}$ and accuracy
3: Base trust score: $\text{trust}_{j}=(1-\bar{u})\cdot(w_{a}\cdot\text{acc}+(1-w_{a}))$
@@
5:  Apply penalty: $\text{trust}_{j}\leftarrow\text{trust}_{j}\cdot\exp(-(\bar{u}-\tau_{u}))$
6:end if
7:return $\text{trust}_{j}$
@@ V-D Adaptive Trust Thresholds
Trust thresholds should adapt as training progresses. Early in training, all models exhibit high uncertainty simply because they have not converged. Applying strict thresholds would reject all peers, preventing collaboration. As training proceeds, models accumulate evidence on their respective distributions and uncertainty decreases for in-distribution data. Following principles from our prior work [rangwala2025sketchguard], we implement adaptive tightening:
@@
where $t$ is the current round and $T$ is the total number of rounds. This starts lenient (accepting peers despite high uncertainty) and gradually tightens (requiring lower uncertainty), matching the expected reduction in epistemic uncertainty as models converge on their respective data distributions. Optional exponential moving average smoothing reduces noise in trust scores across rounds, providing temporal stability to compatibility assessments.
@@ V-E Trust-Aware Aggregation
0: Validation set $\mathcal{D}_{i}^{\text{val}}$, self-weight $\alpha$, round $t$
1: Compute adaptive threshold $\tau_{\min}^{(t)}$
2: Evaluate each neighbor, retaining only those with $\text{trust}_{j}\geq\tau_{\min}$
3:if no peers pass threshold then
4:  return $\boldsymbol{\theta}_{i}$ {Keep local model}
5:end if
6: Normalize trust scores: $w_{j}=\text{trust}_{j}/\sum_{k}\text{trust}_{k}$
7: Aggregate trusted peers: $\boldsymbol{\theta}_{\text{peers}}=\sum_{j}w_{j}\cdot\boldsymbol{\theta}_{j}$
8: Personalize: $\boldsymbol{\theta}_{i}^{\text{new}}=\alpha\cdot\boldsymbol{\theta}_{i}+(1-\alpha)\cdot\boldsymbol{\theta}_{\text{peers}}$
9:return $\boldsymbol{\theta}_{i}^{\text{new}}$
@@ VI Implementation
This section describes the Murmura implementation, detailing the software architecture, key components, and practical considerations for deployment on resource-constrained wearable devices.
@@ VI-A Software Architecture and Design Principles
Murmura is implemented in Python 3.12 using PyTorch [paszke2019pytorch] as the deep learning backend. The framework adopts a modular, layered architecture that cleanly separates concerns and facilitates extensibility. The codebase organization reflects the conceptual model presented in Section IV, with each major component residing in its own package.
@@ VI-B Evidential Deep Learning Components
The evidential modeling capability requires two key components: an evidential output layer and a specialized loss function. The evidential output layer replaces the standard softmax head with a module that produces Dirichlet concentration parameters. This layer consists of a linear transformation from feature space to logit space, followed by an exponential activation ensuring non-negative evidence values, and finally the addition of one to obtain valid Dirichlet parameters greater than zero.
This modification is remarkably lightweight, adding only a single linear layer and activation function beyond a standard classifier. Evidential models compute uncertainty in a single forward pass, avoiding the computational overhead of ensemble methods (which require 5+ model evaluations) or Monte Carlo dropout (requiring dozens of stochastic passes) while providing richer information than point estimates. The computational overhead is empirically less than 1% additional inference time compared to softmax classification.
@@ VI-C Trust-Aware Aggregator Implementation
The evidential trust aggregator implements Algorithm 3 with practical enhancements for stability and efficiency. Trust scores are computed through cross-evaluation with evaluation limited to 100 validation samples to bound computational cost per neighbor. For networks with high connectivity, this limitation ensures that compatibility assessment remains tractable while providing sufficient statistical confidence in uncertainty estimates.
The adaptive trust threshold implements tightening as described in Section V. Optional exponential moving average (EMA) smoothing provides temporal stability to trust assessments. Per-neighbor trust histories are maintained, and new trust scores are blended with historical values using momentum parameter $\gamma\in(0,1)$. Higher momentum values ($\gamma>0.7$) enable rapid adaptation to changing peer behavior. Lower values ($\gamma<0.5$) provide stability against noisy evaluations, beneficial in high-heterogeneity settings where trust scores may fluctuate due to legitimate distribution differences rather than reliability changes.
The self-weight parameter $\alpha$ controlling personalization admits tuning based on expected heterogeneity. In homogeneous settings where all nodes have similar distributions, lower $\alpha$ values ($\alpha<0.3$) enable aggressive collaboration. In highly heterogeneous settings, higher values ($\alpha>0.7$) preserve personalization by limiting peer influence. Our experiments suggest $\alpha=0.5$ provides a reasonable default balancing collaboration and personalization across varied heterogeneity levels.
@@ VI-D Wearable Sensor Dataset Integration
The framework provides adapters for three established wearable datasets (UCI HAR [anguita2013public], PAMAP2 [reiss2012introducing], PPG-DaLiA [reiss2019deep]), each supporting multiple partitioning strategies. Critically, adapters support both synthetic Dirichlet partitioning (for controlled heterogeneity) and natural partitioning by subject ID (for realistic distribution patterns). This dual capability enables validation that findings from controlled experiments generalize to deployment-like conditions where heterogeneity arises from genuine user diversity.

### v2

@@ I Introduction
Wearable Internet of Things (IoT) devices such as smartwatches, fitness trackers, and medical monitors continuously collect sensor data for applications ranging from human activity recognition to early disease detection [2]. Training machine learning models on such data traditionally requires centralized collection, raising privacy concerns and imposing substantial communication and compute overheads on bandwidth-constrained edge networks.
Federated Learning (FL) addresses these concerns by enabling collaborative model training while keeping data on user devices [13]. However, standard centralized FL approaches create a single point of failure and scalability bottleneck problematic for IoT deployments spanning thousands of heterogeneous devices. Decentralized Federated Learning (DFL) eliminates this bottleneck through peer-to-peer architectures where devices exchange and aggregate model updates directly [10], improving fault tolerance and reducing communication costs.
@@ I-B The Peer Compatibility Problem
Existing personalized FL methods typically assume centralized coordination where a server decides how to cluster or weight clients [20]. In fully decentralized settings, each node must make local decisions about peer compatibility without central oversight.
@@ I-C Our Approach: From Uncertainty to Trust
We address this by leveraging uncertainty quantification. When a model evaluates data, we need to separate two fundamentally different types of uncertainty: uncertainty due to lack of knowledge about the data distribution (which signals incompatibility), and uncertainty due to inherent ambiguity in the data itself (which does not). Recent advances in evidential deep learning (EDL) [18] enable exactly this decomposition by having neural networks output not just predictions, but also quantified measures of these two uncertainty types.
@@ I-D Contributions
  • We introduce the first framework combining evidential uncertainty quantification with trust-aware personalization in DFL, enabling nodes to maintain personalized models through selective collaboration in non-IID environments.
@@ II-A Decentralized Federated Learning
FL was introduced to enable collaborative model training while keeping data on local devices [13]. The original approach uses a central server to coordinate training rounds: clients train locally, send updates to the server, which aggregates them and broadcasts the result. While this protects data privacy, it creates centralization challenges.
Decentralized FL removes the need for a central server entirely. Instead, nodes form a network and exchange model updates directly with their neighbors. This peer-to-peer approach improves fault tolerance and can reduce communication costs by leveraging local network structure. Researchers have established both theoretical convergence guarantees [10, 9] and practical benefits including faster training on bandwidth-limited networks [12]. For IoT networks where devices may join and leave dynamically, gossip-based communication provides particularly natural patterns for information exchange [8].
II-B Personalized Decentralized Federated Learning
When devices have heterogeneous data distributions, forcing everyone to share a single global model often produces mediocre results. Personalized FL addresses this by letting each device maintain a model adapted to its own data while still benefiting from collaboration. Most personalized FL methods assume centralized control where a server decides how to group similar clients or weight their contributions [19, 11].
In decentralized settings without a server, personalization becomes harder as nodes must autonomously decide which peers to collaborate with. Recent works have explored peer selection mechanisms based on various similarity measures. BALANCE [6] uses Euclidean distance between model parameters with adaptive thresholds to filter dissimilar peers, originally designed for Byzantine resilience but effectively enabling personalization through distance-based peer selection. Sketchguard [15] employs L2 distance on compressed model sketches to identify compatible peers while reducing communication costs. UBAR [7] combines distance-based screening with loss-based evaluation in a two-stage filtering process. DisPFL [3] and PFedDST [5] use gradient cosine similarity to assess peer compatibility for personalized aggregation. These approaches demonstrate that selective collaboration can improve personalization in decentralized settings.
@@ II-C Evidential Deep Learning
Standard neural network classifiers output probabilities for each class but cannot express confidence in these predictions. EDL [18] addresses this by having networks output evidence values that parameterize a Dirichlet distribution over probability vectors. This enables explicit quantification of epistemic uncertainty (lack of knowledge due to insufficient training or distribution shift) and aleatoric uncertainty (inherent data ambiguity). Epistemic uncertainty directly indicates whether a model has learned relevant patterns for given data, making it particularly useful for identifying compatible peers in heterogeneous FL settings. While EDL has been used in various FL contexts such as selecting reliable models [22] or weighting aggregation in centralized settings [21], no prior work has leveraged it for compatibility-based peer selection in fully decentralized personalized FL under non-IID data conditions.
@@ II-D Positioning and Research Gap
Table I compares Murmura with related approaches across key dimensions. Existing methods fall into two categories, each with fundamental limitations for our setting.
Decentralized personalized FL methods [6, 15, 7, 3, 5] use heuristic similarity measures for peer selection. Distance-based methods treat all parameter differences equally regardless of whether they reflect distributional mismatch or legitimate model diversity. Gradient cosine similarity conflates training quality with distributional compatibility: two peers might have dissimilar gradients because one is well-trained on different data (incompatible but high-quality) or poorly trained on similar data (compatible but low-quality). These scenarios require opposite aggregation decisions but cannot be distinguished without understanding why models differ.
Uncertainty-aware aggregation methods face different limitations. Centralized approaches like [21, 19, 11] rely on a server to evaluate uncertainty across all clients and make global decisions, unavailable when nodes operate autonomously. Prediction-level uncertainty methods cannot decompose epistemic (lack of knowledge) from aleatoric (inherent ambiguity) uncertainty. This decomposition is critical: a peer model showing high prediction uncertainty could indicate either incompatible training data or legitimately challenging samples. Without epistemic uncertainty quantification, nodes cannot distinguish these cases.
Murmura addresses both limitations by enabling autonomous compatibility assessment through epistemic uncertainty. High epistemic uncertainty during cross-evaluation directly reveals distributional mismatch regardless of prediction accuracy or gradient characteristics. The key novelty is applying evidential uncertainty decomposition to enable principled peer selection in fully decentralized personalized FL.
@@
[19] Central
[11] Central
[6] L2 distance
[15] L2 distance
[7] Distance+Loss
[3] Gradient sim.
[5] Gradient sim.
[21] Central
Murmura Epistemic
@@ III-D Design Objectives
Murmura is designed to satisfy four objectives:
@@ IV Murmura Framework
This section presents the Murmura framework, emphasizing the architectural decisions that enable evaluation of trust-aware personalization under controlled heterogeneity.
@@ IV-A Design Principles and Architecture
Murmura achieves these goals through modular separation of concerns, as illustrated in Figure 1. The architecture comprises three layers. The Core layer separates network orchestration (managing training rounds and communication) from node-level decision making (local training and aggregation). This separation is essential: it allows individual nodes to implement different aggregation strategies while maintaining identical communication patterns, isolating the impact of personalization mechanisms from topology effects.
Figure 1: Murmura framework architecture showing the three-layer design. The Configuration Layer provides YAML/JSON-based experiment specification. The Core Components layer contains the Network Orchestrator and Node, along with Topology and Model definitions. The Pluggable Modules layer enables flexible integration of aggregation strategies, data partitioning schemes and datasets.
The Pluggable Modules layer decouples topology generation from aggregation logic. This enables systematic study of how personalization strategies perform under different network structures (ring, fully-connected, Erdős-Rényi [4], k-regular) without conflating topology and algorithm effects—a common limitation in prior work where algorithms are evaluated on fixed topologies.
@@ V Evidential Trust-Aware Personalization
This section explains how Murmura uses EDL to assess peer compatibility and enable personalized model aggregation in heterogeneous data environments.
@@ V-B Training Evidential Models
Evidential models require a specialized loss function that encourages accurate predictions while preventing the model from simply claiming high confidence everywhere. Following Sensoy et al. [18], we use:
@@
where $\mathbf{y}=[y_{1},\ldots,y_{K}]$ is the one-hot encoded true label, $\hat{p}_{k}=\alpha_{k}/S$ is the expected probability for class $k$, and $\tilde{\boldsymbol{\alpha}}=\mathbf{y}+(1-\mathbf{y})\odot\boldsymbol{\alpha}$ removes evidence from the true class (where $\odot$ denotes element-wise multiplication). The first term is a squared error measuring prediction accuracy, while the KL divergence term prevents accumulating evidence for incorrect classes. The coefficient $\lambda_{t}$ controls the strength of the KL regularization and follows a linear annealing schedule:
$\lambda_{t}=\lambda_{\max}\cdot\min\left(1,\frac{t}{T_{\text{anneal}}}\right)$ (7)
where $t$ is the current training round, $T_{\text{anneal}}$ is the number of rounds over which to anneal (typically set to half the total training rounds), and $\lambda_{\max}$ is the maximum regularization weight. This linear annealing allows initial learning of accurate predictions before regularizing uncertainty estimates, preventing premature commitment to low-confidence predictions.
@@ V-C Computing Compatibility from Uncertainty
2: Compute mean epistemic uncertainty $\bar{u}$ and accuracy $a$
3: Base trust score: $s_{\text{base}}=(1-\bar{u})\cdot(w_{a}\cdot a+(1-w_{a}))$
@@
5:  Apply penalty: $s_{\text{final}}=s_{\text{base}}\cdot\exp(-(\bar{u}-\tau_{u}))$
6:else
7:  $s_{\text{final}}=s_{\text{base}}$
8:end if
9:return $s_{\text{final}}$
@@ V-D Adaptive Trust Thresholds
Trust thresholds should adapt as training progresses. Early in training, all models exhibit high uncertainty simply because they have not converged. Applying strict thresholds would reject all peers, preventing collaboration. As training proceeds, models accumulate evidence on their respective distributions and uncertainty decreases for in-distribution data. Following principles from [15], we implement adaptive tightening:
@@
where $t$ is the current round, $T$ is the total number of rounds, $\tau_{\min}^{(0)}$ is the initial minimum trust threshold, $\gamma_{\tau}\in[0,1]$ controls the maximum tightening amount, and $\kappa>0$ controls the tightening rate. This schedule starts lenient (accepting peers despite high uncertainty) and gradually tightens (requiring lower uncertainty), matching the expected reduction in epistemic uncertainty as models converge on their respective data distributions.
@@ V-E Trust-Aware Aggregation
0: Validation set $\mathcal{D}_{i}^{\text{val}}$, self-weight $\omega$, round $t$
1: Compute adaptive threshold $\tau_{\min}^{(t)}$ using Equation (6)
2: Evaluate each neighbor using Algorithm 2, obtaining trust scores $\{s_{j}\}$
3: Retain only peers with $s_{j}\geq\tau_{\min}^{(t)}$
4:if no peers pass threshold then
5:  return $\boldsymbol{\theta}_{i}$ {Keep local model}
6:end if
7: Normalize trust scores: $w_{j}=s_{j}/\sum_{k}s_{k}$ for retained peers
8: Aggregate trusted peers: $\boldsymbol{\theta}_{\text{peers}}=\sum_{j}w_{j}\cdot\boldsymbol{\theta}_{j}$
9: Personalize: $\boldsymbol{\theta}_{i}^{\text{new}}=\omega\cdot\boldsymbol{\theta}_{i}+(1-\omega)\cdot\boldsymbol{\theta}_{\text{peers}}$
10:return $\boldsymbol{\theta}_{i}^{\text{new}}$
@@ VI Implementation
This section describes the Murmura [^1] The complete implementation of Murmura is available as open-source software at https://github.com/Cloudslab/murmura. implementation, detailing the software architecture, key components, and practical considerations for deployment on resource-constrained wearable devices.
@@ VI-A Software Architecture and Design Principles
Murmura is implemented in Python 3.12 using PyTorch [14] as the deep learning backend. The framework adopts a modular, layered architecture that cleanly separates concerns and facilitates extensibility. The codebase organization reflects the conceptual model presented in Section IV, with each major component residing in its own package.
@@ VI-B Evidential Deep Learning Components
The evidential modeling capability requires two key components: an evidential output layer and a specialized loss function. The evidential output layer replaces the standard softmax head with a module that produces Dirichlet concentration parameters. This layer consists of a linear transformation from feature space to logit space, followed by an exponential activation ensuring non-negative evidence values, and finally the addition of one to obtain valid Dirichlet parameters. This modification is remarkably lightweight, adding only a single linear layer and activation function beyond a standard classifier. Evidential models compute uncertainty in a single forward pass, avoiding the computational overhead of ensemble methods or Monte Carlo dropout while providing richer information than point estimates. The computational overhead is empirically less than 1% additional inference time compared to softmax classification.
@@ VI-C Trust-Aware Aggregator Implementation
The evidential trust aggregator implements Algorithm 3 with practical enhancements for stability and efficiency. Trust scores are computed through cross-evaluation, with evaluation limited to 100 validation samples to bound computational cost per neighbor. For networks with high connectivity, this limitation ensures that compatibility assessment remains tractable while providing sufficient statistical confidence in uncertainty estimates. The adaptive trust threshold implements tightening as described in Section V using the exponential decay schedule with configurable parameters $\gamma_{\tau}$ and $\kappa$ that control the tightening behavior. Two key parameters control the personalization behavior. First, the self-weight parameter $\omega$ in Algorithm 3 determines how much influence the aggregated peer models have versus the local model. Lower $\omega$ values ($\omega<0.3$) enable aggressive collaboration in homogeneous settings where all nodes have similar distributions. Higher values ($\omega>0.7$) preserve personalization by limiting peer influence in highly heterogeneous settings. Second, the accuracy weight $w_{a}$ in Algorithm 2 balances the importance of prediction accuracy versus uncertainty characteristics when computing trust scores for peer compatibility evaluation.
@@ VI-D Wearable Sensor Dataset Integration
The framework provides adapters for three established wearable datasets (UCI HAR [1], PAMAP2 [17], PPG-DaLiA [16]), each supporting multiple partitioning strategies. Critically, adapters support both synthetic Dirichlet partitioning (for controlled heterogeneity) and natural partitioning by subject ID (for realistic distribution patterns). This dual capability enables validation that findings from controlled experiments generalize to deployment-like conditions where heterogeneity arises from genuine user diversity.